OpenStudy (anonymous):

MEDAL FOR HELP AYYYE LMAO Find 4 rational numbers such that the product of the second,third, and fourth number = -63. The second number is 2 more than the first number, the third is 5 less than twice the second, and the fourth is six less than the first.

myininaya (myininaya):

Ok so let's say we have the rational numbers a,b,c, and d We are given: The product of b,c and d is -63. b is 2 more than a. c is 5 less than twice the b. d is 6 less than a . Do you know how to translate the givens into mathematical equations.

OpenStudy (anonymous):

The answer is \[{ -3, -1, -7, -9}\] \[{2.5, 4.5, 4, -3.5}\] \[{5, 7, 9, -1}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so i got the first is x, the second is x + 2, the third is 2x-1, and the fourth x-6

OpenStudy (anonymous):

idk what to do after that

myininaya (myininaya):

how do you know they are in that form?

OpenStudy (anonymous):

What do you mean?

myininaya (myininaya):

ok so i got the first is x, the second is x + 2, the third is 2x-1, and the fourth x-6 how did you find this out?

OpenStudy (anonymous):

ok if the second is 2 more than the first (x) then its x +2

OpenStudy (anonymous):

then the third is 2(x+2) -5

OpenStudy (anonymous):

=2x-1

myininaya (myininaya):

So are you saying that it is given the the first, second, third, and fourth numbers they are talking is x,x+2,2x-1,x-6?

OpenStudy (anonymous):

the 4th is six less than the first (x) -6

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

yeah its given :P

OpenStudy (anonymous):

do u then set it up into a polynomial? (x+2)(2x-1)(x-6) = -63 ?

myininaya (myininaya):

ok so... Ok so let's say we have the rational numbers a,b,c, and d We are given: The product of b,c and d is -63. b is 2 more than a. c is 5 less than twice the b. d is 6 less than a . Do you know how to translate the givens into mathematical equations. instead of me writing a,b,c, and d then we can write x,x+2,2x-1,x-6 So you are write that first sentence translates to: (x+2)(2x-1)(x-6)=-63 So let's look at solving: (x+2)(2x-1)(x-6)=-63

OpenStudy (anonymous):

ok

myininaya (myininaya):

I just copied paste some of that so I didn't have to look back up

OpenStudy (anonymous):

thats cool

myininaya (myininaya):

This is a cubic equation

OpenStudy (anonymous):

so it simplifies to \[2x^3-11x^2-8x+75 = 0\]

myininaya (myininaya):

let me check

OpenStudy (anonymous):

i may have miscalculated but thats what I got

myininaya (myininaya):

two terms are off

myininaya (myininaya):

but you did good on the first and last term on the left hand side

myininaya (myininaya):

\[(x+2)(2x-1)(x-6)=-63 \\ (x+2)(2x-1)(x-6)+63=0 \\ [(x+2)(x-6)](2x-1)+63=0\] let's look at multiplying first (x+2) and (x-6)

myininaya (myininaya):

\[(x^2-4x-12)(2x-1)+63=0\] see if you can (x^2-4x-12) and (2x-1)

OpenStudy (anonymous):

\[2x^3-x^2-8x^2+4x-24x+12\]

OpenStudy (anonymous):

?

myininaya (myininaya):

that actually looks lovely and then we have +63 to tact on

OpenStudy (anonymous):

\[2x^3-9x^2-20x+75\]

myininaya (myininaya):

\[2x^3-9x^2-20x+75=0\]

myininaya (myininaya):

now let's see if we can apply the rational root theorem

OpenStudy (anonymous):

ooooooohhhh

OpenStudy (anonymous):

ok i got it

myininaya (myininaya):

that is we look at the factors of 75 over the factors of 2 \[\pm \frac{75}{2}, \pm\frac{25}{2}, \pm \frac{15}{2}, \pm \frac{5}{2}, \pm \frac{3}{2}, \pm \frac{1}{2}, \pm \frac{75}{1}, \pm \frac{25}{2}, \pm \frac{15}{1}, \pm \frac{5}{1}, \pm \frac{3}{1}, \pm \frac{1}{1}\] that is a lot of numbers to check :p

myininaya (myininaya):

but you got it you say?

OpenStudy (anonymous):

Yeah our teacher allows graphing calculators

myininaya (myininaya):

lol oh that makes it a tad easier

OpenStudy (anonymous):

yeaaahh XD thanks for the help!

myininaya (myininaya):

np